OpenStudy (anonymous):

Use substitution method. Answer as an ordered pair. 2x-y=10 3x-2y=8

sam (.sam.):

\[2x-y=10...(1)\] \[3x-2y=8...(2)\] Isolate y by adding y and subtracting 10 to both sides \[2x-y\color{red}{+y-10}=10\color{red}{-10+y}\] we get \[2x-10=y...(3)\] now substitute (3) into (2) \[3x-2y=8\] \[3x-2(2x-10)=8\] solve for x